OpenStudy (anonymous):

solve : 1. ) 2/43=j/13

OpenStudy (mayankdevnani):

\[\huge \frac{2}{43}=\frac{j}{13}\] do cross multiply, \[\huge 26=43j\] \[\huge j=\frac{26}{43}\]